# Do Automatic Factuality Metrics Measure Factuality? A Critical Evaluation

### v3

@@ 4.2 Measuring Metric Sensitivities to Controlled Manipulations
Results
We present our results in Figure 4.2. Specialized models MiniCheck and AlignScore, and prompt-based ChatGPT-DA show the strongest positive score shifts following factual corrections, as one would hope for: Corrected, more factual summaries are rated higher than inconsistent ones. SummaC-Conv and UniEval also show positive, albeit smaller, differences. QuestEval shows no change (on average) for corrected summaries.
@@ Results
Most metrics end up showing significant score increases when a random source sentence is inserted into an inconsistent summary—often surpassing or matching the gains from factual corrections. [^2] Such a sentence will itself be “consistent”, but should probably not shift the overall consistency score upwards given that this does not correct any existing inconsistencies, as it strictly adds more content. QuestEval and SummaC-Conv exhibit the largest increases in this dimension, while AlignScore and UniEval show comparable improvements to corrected scores. In contrast, MiniCheck and ChatGPT-DA exhibit positive shifts but smaller than those observed for corrected summaries.
Another clear pattern emerges across all metrics, especially with MiniCheck, UniEval, and ChatGPT-DA show a decrease in scores when meaning-preserving negation terms are introduced. These negations—added through simple rephrasing without changing the original meaning (e.g., changing “The author mistakenly believed their ACT test was on a certain day” to “The author did not realize their ACT test was not on the day they thought”)—suggest that these models are particularly sensitive to negation.
SummaC and UniEval are more sensitive to benign modifications than to corrections. In contrast, AlignScore and MiniCheck exhibit greater sensitivity across a range of benign modifications. While ChatGPT is sensitive to negated sentences, it tends to be more robust to other types of modifications.
Taken together, these results demonstrate the variability of factual consistency metrics with respect to spurious and meaningful edits to outputs. And given that at least some of the metrics appear to be responsive to inconsequential edits, our results imply that we might be able to game metrics, i.e., systematically inflate scores without attendant changes in factual consistency. [^3] One could argue that such manipulations—adversarial ones especially—result in “out of distribution” inputs, and so we should have no expectation of how models will perform. But usually factual consistency metrics are touted (implicitly) as measuring consistency between arbitrary document and summary candidate pairs.
@@ 5 Can we game factuality metrics?
Constant Phrase 2
The summary entails the information the document discusses. 0.93
Manipulated Summary
@@
Table 2: Qualitative (cherry-picked) samples of original and manipulated summaries with corresponding metric scores for AlignScore and MiniCheck. For comprehensiveness, we report quantitative aggregated results in Figure 5.1, and we provide more examples in Appendix 7.
Figure 5.1: Pairwise score differences across summaries manipulated using four distinct strategies: Adding constant strings (top phrase and assertion phrase) and appending them to summaries (summ + top, and summ + assertion. The results reveal that NLI and bespoke model metrics are particularly vulnerable to gaming, with significant score inflation observed under these manipulations.
Strings for summary manipulation We attempt to identify a set of strings that artificially inflate metric scores. To do so, we first identify the top 20th percentile of summaries ranked by each factuality metric, reasoning that these high-scoring summaries may contain patterns that raise scores, independent of context. To this end, we compute TF-IDF scores for all bigrams in these summaries, identifying those disproportionately present in summaries deemed “factual” by each metric.
We select from these the top 100 bigrams, aggregating results across metrics. This set includes, e.g., “the document” and “document discusses”. We adopt the constant phrase “the document discusses” (top phrase) into all documents and observe whether metrics are responsive to this. We consider one additional phrase: such as “The summary entails information in the document.” (assertion phrase) which explicitly asserts factual consistency. The wording of this phrase varies slightly across metrics to align with their specific methods for evaluating factual consistency. The complete list of these phrases for each metric is provided in the Appendix D.1. Finally, we append both phrases to the corresponding summaries summ (+top) and summ (+assertion).
Our goal here is to determine whether the appended text unduly influences metrics, potentially resulting in inflated scores that do not accurately reflect the factual quality of the original content.
@@
Metric sensitivity Figure 5.1 reports the effects of our gaming strategies on metric scores. Notably, the constant phrases boosts scores by >>>0.2 points (absolute) for NLI-based SummaC-Conv, as well as specialized models UniEval, AlignScore, and MiniCheck. This is surprising, as these phrases are not valid summaries and do not contain any factual content (note that top phrase is an incomplete sentence). Adding constant phrases as suffixes to summaries increased scores by 0.1–0.15 points; this is comparable to the gains realized following factual corrections, as reported in Section 4.2. In fact, SummaC-Conv shows no score increase for corrected summaries, suggesting an under-sensitivity to actual changes in factuality.
To further contextualize these numbers, we analyze metric score differences between summaries generated by larger models (e.g., GPT-4, Gemini) and smaller models (e.g., Llama-7B, Mistral-7B, Falcon-7B, BART). Larger models generally produce more consistent summaries (Tang et al., 2022; Goyal et al., 2022), providing a baseline for assessing gaming strategies. A natural question, then, is whether “gaming” achieves score improvements comparable to those that have resulted from advances in model size and complexity.
We analyze a subset of 115 document-summary pairs from GenAudit (Krishna et al., 2024) and Aggrefact-old (Tang et al., 2022), where model information is available. Our findings reveal that gaming with constant phrases leads to metric score improvements 3x to 8x greater than those achieved through genuine improvements in generated summaries, particularly for NLI and specialized metrics. For instance, gaming increases SummaC scores by nearly 0.2 points, while model improvements have minimal impact. Similarly, adding suffixes to summaries achieves gains comparable to those won with better summarization models. (Full results in Appendix D.1.)
QuestEval is mostly unaffected by these manipulations. ChatGPT-DA shows a negative trend: Scores drop when summaries are “gamed”. Further analysis suggests that the QA-based metric generates questions based on innocuous phrases and entity repetitions that introduce noise, leading to unchanged or lower scores. ChatGPT-DA’s performance decline likely results from its strict adherence to the prompts. A qualitative analysis, where the model was prompted to explain its scoring, revealed that in several instances, it explicitly identified the gaming phrase as misleading. However, we observe that ChatGPT-DA’s scores can be inflated by replacing summaries with a disclaimer acknowledging potential errors in interpreting the source (see Appendix D.2).
@@
Overall, these results suggest that adding fixed phrases to summaries can boost metric scores—for most finetuned model metrics—at levels comparable to, or even exceeding, presumably genuine improvements due to advances in summarization models themselves.
Filler phrases and metric performance
The innocuous, non-factual phrases used in our experiments above constitute “filler” phrases, which introduce no substantive content. Since most factuality metrics primarily assess factual precision, we hypothesize that they may struggle to differentiate between informative content and irrelevant filler text leading to a weakness that could be exploited. Our experiments above provide some evidence of this and we explore it further in this section. First, in addition to “gaming prefixes” we consider two other types of constant strings that are neutral phrases: They are independent of our analysis of top-scoring bigrams and do not allude to the consistency of summaries with respect to the document. We measure the pairwise difference in metric scores when these phrases are appended as a point of reference.
We consider two neutral phrase types: a “baseline” phrase, consisting of two generic phrases that don’t mention the summary or source, and a “qualifier” phrase, suggesting the summary may have multiple interpretations (see Table 6 for phrases). This contrasts with our “assertion” phrase, which directly affirms the summary’s faithfulness. These filler phrases should ideally not affect consistency scores, as they don’t alter factual content. We also compare how metrics respond to these phrases versus “gaming” phrases to determine if only targeted fillers increase scores.
As shown in Appendix D.2, all metrics assign lower scores to summaries with a baseline phrase, indicating that not all irrelevant sentences inflate scores. However, qualifier phrases have similar effects to gaming prefixes on MiniCheck and UniEval, suggesting that neutral phrases referencing the summary may boost scores for these metrics. For SummaC and AlignScore, only specific gaming prefixes increase scores, while other phrases reduce them. Interestingly, ChatGPT, which usually shows a negative trend, exhibits inflated scores when the summary is replaced with a neutral qualifier phrase acknowledging potential errors in interpreting the source.
Although human evaluators may accept such neutral phrases as consistent, they likely assess models more comprehensively, identifying deeper inconsistencies. In contrast, metrics are not sophisticated enough to capture these nuances and can be influenced by irrelevant information, inflating scores independent of factual consistency.
6 Discussion
Distilling salient information from large volumes of unstructured text—i.e., summarization—is a compelling practical use case for LLMs. As they have grown increasingly competent at this task, it has become concomitantly difficult to evaluate their performance, which requires assessing often subtle but potentially critical factual inconsistencies between inputs and generated summaries. Safe deployment of, and continued progress on, such approaches to summarization requires automatic metrics to assess factual consistency automatically. Many such approaches have been proposed in the literature, but there has been little effort to stress-test them, which is the goal of the present work.
We found first that a simple MLP fine-tuned over superficial features and summaries can realize performance competitive with SOTA factual consistency metrics. We then showed that while such metrics are generally responsive to manual changes affecting factual consistency, they are also often sensitive to spurious transformations which should not affect scores. Motivated by these findings, we then showed that factual consistency metrics are often gameable, i.e., in many cases we can add fixed phrases as suffixes into output summaries or use them standalone to reliably inflate “factuality” scores beyond gains observed via improvements in model size and complexity.
@@ 6 Discussion
First, in Section 3, we demonstrated that a model trained solely on heuristic indicators of factual consistency can achieve performance on par with state-of-the-art (SOTA) approaches, though this was a supervised model specifically for consistency detection. Despite our training data being sourced from entirely different datasets and domains than those used for evaluation, it is important to highlight that the dataset was manually annotated. However, we cannot yet conclude how a shallow model would perform if trained on heuristic-based predictors of synthetically generated summaries—this would be most similar to the data that other (specialized) metrics have been trained with. We emphasize, however, that our goal with these experiments was not to introduce a new metric, but rather to investigate how much current metrics improve upon basic textual attributes.
Another limitation here is that one could quibble with just how “shallow” these predictors are. Our view is that even the more involved features we use for this—e.g., semantic similarity based on BERT embeddings and entity overlap—should not be taken as reliable measures of factual consistency, but empirically they do correlate with manual annotations of factual consistency (see Appendix B). We can only speculate regarding the underlying causal structure based on these results.
Our subsequent interventions in Section 4.2 established that in some cases metrics are sensitive to extraneous transformations (e.g., re-phrasings). However this was not observed uniformly, and we performed such transformations using LLMs, which may have introduced legitimate factual inconsistencies (though we believe this to be rare). A limitation to our “gaming” experiment is that it involved a manual process; however, this does not affect the main result, which is that there exist strings that can be appended to outputs in such a way that consistently boosts scores.
A further limitation to this (and all manipulation experiments) is that they produce outputs which may be viewed as “out of distribution”; but generally the purpose of summary factuality metrics is to be able to apply them to arbitrary model outputs and corresponding references, so it is not entirely clear what “in distribution” means under this assumption.
Despite these limitations, we believe that—taken together—our results suggest that one should be careful about their interpretation of automatic factual consistency metrics. And we hope these results inform research on summarization moving forward.
Appendix A Experimental Setup
A.1 Benchmark Datasets
We list the datasets used in the benchmarks and specify which split (train/test) they are included in Table 3.
@@ A.1 Benchmark Datasets
LLM-Aggrefact (Tang et al., 2024a) TofuEval-MeetB (Tang et al., 2024b),
ExpertQA (Malaviya et al., 2023)
Lfqa (Chen et al., 2023)
RAGTruth (Wu et al., 2023)
FactCheck-GPT (Wang et al., 2023b)
Wice (Kamoi et al., 2023a)
TofuEval-MediaS (Tang et al., 2024b)
ClaimVerify (Liu et al., 2023)
Reveal (Jacovi et al., 2024)
Aggrefact (Tang et al., 2022) FRANK (Pagnoni et al., 2021)
Polytope (Huang et al., 2020)
FactCC (Kryscinski et al., 2020)
Goyal21 (Goyal and Durrett, 2021)
Wang20 (Wang et al., 2020)
Cao22 (Cao et al., 2022)
XSumFaith (Maynez et al., 2020)
CLIFF (Cao and Wang, 2021)
SummEval (Fabbri et al., 2021a)
FacEval (Wang et al., 2022) SAMSUM (Gliwa et al., 2019b) -
LLM-Dialogue (Ramprasad et al., 2024a) SAMSUM DialogSum (Chen et al., 2021b)
Genaudit (Krishna et al., 2024) - REDDIT,
ACIBENCH (Yim et al., 2023),
XSUM (Narayan et al., 2018)
Table 3:
Appendix B Metric correlation with shallow features
Figure B.1: Spearman rank correlations between each automatic evaluation metric and individual shallow features. Some of these correlations are moderate, suggesting that in aggregate superficial features may carry much of the signal being measured by automated measures of factual consistency.

### v4

@@ 4.1 Measuring Metric Sensitivities to Controlled Manipulation
Results.
@@
We present results in Figure 3. Notably, QuestEval shows no meaningful improvement in response to factual corrections, raising concerns about its ability to detect genuine consistency improvements. Additionally, most metrics, except QuestEval and ChatGPT-DA, assign disproportionately high scores when random, yet consistent, source sentences are appended to inconsistent summaries. [^3] Such a sentence will itself be “consistent”, but should probably not shift the overall consistency score to be higher than the corrected summary given that this does not correct any existing inconsistencies, as it strictly adds content. In many cases, these spurious additions yield score gains comparable to, or greater than, those resulting from actual factual corrections.
We also observe that MiniCheck and UniEval are highly sensitive to meaning-preserving negation. Simple rephrasings that retain the original meaning—for example, changing “The author mistakenly believed their ACT test was on a certain day” to “The author did not realize their ACT test was not on the day they thought”, often lead to reduced scores. While ChatGPT-DA exhibits some sensitivity to negation, it remains relatively robust overall, consistently distinguishing between true factual improvements and irrelevant textual changes.
4.2 Reliance on Source Context Versus Parametric Knowledge in GPT-Based Scoring
Figure 4: GPT-based consistency evaluation is influenced by parametric knowledge. Left: The score gap between supported and unsupported summaries narrows sharply when references are counterfactual but summaries are factually accurate (p < 0.001). Right: The rate of cases where unsupported summaries are scored higher than supported ones rises from 0.2% to 3.1% when references contradict GPT’s world knowledge while summaries remain factually correct.
Prompt-based approaches to factual consistency evaluation have recently gained traction (luo2023chatgpt; wang2023chatgpt). Unlike traditional metrics trained on source-summary pairs, these methods rely on prompting pre-trained models to evaluate summaries against reference articles for consistency. This raises a critical question: Are their judgments truly grounded in the provided source, or are they influenced by the model’s internal (parametric) knowledge?
We investigate this by evaluating GPT’s consistency judgments under scenarios where the reference text (likely) conflicts with the model’s implicit world knowledge. To ensure that GPT possesses strong prior knowledge about the domain, we use Wikipedia articles, which are heavily represented in LLM pretraining corpora. Specifically, we use the ConflictBank dataset (su2024conflictbank), which comprises factual claims extracted from Wikipedia and corresponding counterfactual variants generated through targeted substitutions. Each counterfactual claim is also paired with a counterfactual reference document—a version of the original article modified to make the false claim appear consistent with the altered reference. This dataset structure permits four experimental conditions: (a) Factual references paired with supported factual summaries; (b) Counterfactual references paired with supported counterfactual summaries; (c) Factual references paired with unsupported counterfactual summaries, and; (d) Counterfactual references paired with unsupported factual summaries.
To test the discriminative ability of GPT to score for consistency, we evaluate against factual references using conditions (a) and (c) and counterfactual references using conditions (b) and (d). Ideally GPT would be equally discriminative in both cases: (a) and (b) should be deemed equally consistent, and (c) and (d) be scored as equally inconsistent. In Figure 4, left, we compare the score differences between supported and unsupported summaries in both settings, namely (a - c) and (b - d), using a paired t-test on the two deltas. The test yields p < 0.001, i.e., the score gap between supported and unsupported summaries is significantly smaller when evaluated against counterfactual references compared to factual ones. This suggests that GPT struggles more to distinguish between supported and unsupported summaries when the reference contradicts its internal knowledge. In particular, when evaluating summaries against references that contradicts its internal knowledge, the model shows reduced sensitivity to whether the summary is actually supported by evidence.
We would also anticipate that the model should score supported summaries higher than unsupported summaries, i.e., b > d. However, if d > b, this might suggest that GPT’s parametric knowledge is overriding the reference-based evaluation, inappropriately favoring (generally) factually accurate content which is unsupported by the input under consideration. We find that GPT incorrectly rates the unsupported (but factual) summary higher than the supported summary in 3% of cases. As a baseline, we also evaluate the proportion of times it scores unsupported outputs higher than supported factual summaries where the input is (generally) factual (c > a). Here, GPT rates unsupported summaries higher only 0.2% of the time. The increase in inconsistent summaries being rated higher ($0.2\%\rightarrow 3\%$) suggests that GPT struggles more with consistency evaluation when references contain misinformation or contradict established facts. This suggests the model’s evaluation reliability could be compromised in real-world scenarios involving contested information, fiction, updated knowledge, or reference materials that contradict GPT’s implicit parametric knowledge.
@@ 5 Can we game factuality metrics?
Figure 5: Pairwise score differences across summaries manipulated using four distinct strategies: Adding constant strings (top phrase and assertion phrase) and appending them to summaries (summ + top, and summ + assertion. The results reveal that NLI and bespoke model metrics are particularly vulnerable to gaming, with significant score inflation observed under these manipulations.
@@
Summary w/ Phrase 1
@@
Table 2: Qualitative (cherry-picked) samples of original and manipulated summaries with corresponding metric scores for AlignScore and MiniCheck. For comprehensiveness, we report quantitative aggregated results in Figure 5, and we provide more examples in Appendix 7.
Strings for summary manipulation. We attempt to identify a set of strings that artificially inflate metric scores by analyzing the top 20th percentile of summaries ranked by each factuality metric, reasoning that these high-scoring summaries may contain patterns that raise scores, independent of context. We then compute TF-IDF scores for bigrams in these high-scoring summaries to detect patterns disproportionately associated with "factual" summaries.
We select from these the top 100 bigrams, aggregating results across metrics. This set includes, e.g., “the document” and “document discusses”. We adopt the constant phrase “the document discusses” (top phrase) into all documents and observe whether metrics are responsive to this. We consider one additional phrase: “The summary entails information in the document.” (assertion phrase) which explicitly asserts factual consistency. [^4] The wording of this phrase varies slightly across metrics to align with their specific methods for evaluating factual consistency. The complete list of these phrases for each metric is provided in the Appendix 12.1 Finally, we append both phrases to the corresponding summaries summ (+top) and summ (+assertion).
@@
Figure 5 reports the effects of our gaming strategies on metric scores. Notably, the constant phrases boosts scores by >>0.2 points (absolute) for NLI-based SummaC-Conv, as well as specialized models UniEval, AlignScore, and MiniCheck. This is surprising, as these phrases are not valid summaries and do not contain any factual content (note that the top phrase is an incomplete sentence). Adding constant phrases as suffixes to summaries increased scores by 0.1–0.15 points; this is comparable to the gains realized following factual corrections (see Section 4.1). Indeed, SummaC-Conv shows no score increase for corrected summaries, suggesting an under-sensitivity to actual changes in factuality.
To contextualize these results, we compare score differences between summaries from larger models (e.g., GPT-4, Gemini) and smaller models (e.g., Llama-7B, Mistral-7B, Falcon-7B, BART). While larger models typically yield more consistent summaries (tang2022understanding; goyal2022news), we find that gaming with constant phrases produces larger score gains—often exceeding those from genuine model improvements, especially for NLI and specialized metrics.
@@
One could argue that such manipulations—adversarial ones especially—result in “out of distribution” inputs, and so we should have no expectation of how models will perform. But usually factual consistency metrics are touted (implicitly) as measuring consistency between arbitrary document and summary candidate pairs. Overall, these results suggest that adding fixed phrases to summaries can boost metric scores—for most finetuned model metrics—at levels comparable to, or even exceeding, presumably genuine improvements due to advances in summarization models themselves.
6 Related Work
Prior work has meta-evaluated factuality metrics, focusing on their sensitivity to specific error types, frequencies, or domains (gabriel-etal-2021-go; chen-etal-2021-factuality-checkers). In contrast, we analyze how metrics respond to factual corrections versus unrelated textual edits, exposing their vulnerability to spurious cues. kamoi-etal-2023-shortcomings also find that QA-based metrics, while effective at summary-level scoring, fail to localize errors and can be outperformed by simple baselines. goyal-durrett-2021-annotating note that no metric consistently outperforms others, though their analysis focuses on error types; we highlight on reliability and potential issues with such metrics.
7 Conclusions and Discussion
Factuality benchmarks should be updated to reflect the evolving nature of errors produced by modern LLMs, which are increasingly subtle and context-dependent. Our analysis in Section 3 reveals that even modern benchmarks include trivial errors that shallow metrics can easily detect. Evaluation should instead prioritize challenging examples that demand deeper semantic reasoning, ensuring that metrics are assessed on their ability to capture nuanced inconsistencies rather than surface-level cues. This may be especially important in complex and high stakes domains like law and medicine.
In this work, we have taken a step in this direction by introducing a shallow feature-based probe to categorize benchmark examples by complexity, enabling more informative evaluations of metric performance. Our analysis in Section 4.1 reveals that many metrics do not reliably distinguish between factual corrections and benign edits, indicating poor calibration to factual severity. We recommend that benchmarks explicitly incorporate graded summary variants reflecting different levels of factual severity. Specifically, each source should be paired with multiple summary edits to test whether metrics appropriately reward factual improvements and remain stable under benign edits. Here we approximated this approach using the dataset from krishna2024genaudit, which includes summaries edited at two levels of factuality alongside benign variants.
In Section 5 we show that most specialized metrics can be gamed with factuality assertion or qualifier phrases that do not alter content, revealing a vulnerability: current metrics often score factuality without accounting for saliency. To address this, we recommend integrating saliency-aware scoring that prioritizes evaluating core content aligned with the source. This may reduce susceptibility to filler-based manipulations and better reflect meaningful consistency.
Prompt-based metrics (using LLMs) show greater robustness to benign edits and gaming, making them strong candidates for factuality evaluation especially in settings where reliability is of concern. However, since they are not explicitly designed for fact-checking, their reliance on pre-trained knowledge must be considered. We recommend the development of benchmarks that specifically evaluate whether LLM-based methods ground their judgments in the source. To assess whether models rely on source content or internal priors, benchmarks might include summaries containing unsupported inferences or source information that contradicts parametric knowledge. In this work we investigated this behavior using synthetic source manipulations from ConflictBank (su2024conflictbank), finding that GPT models often default to their parametric knowledge when evaluating consistency against contradictory references. These results highlight the importance of incorporating misinformation, controversial claims, and fictional content into source-grounded factuality benchmarks
@@ 7 Conclusions and Discussion
Our interventions in Section 4.1 established that in some cases metrics are sensitive to extraneous transformations (e.g., re-phrasings). However this was not observed uniformly, and we performed such transformations using GPT-4, which may have introduced legitimate factual inconsistencies (though we believe this to be rare).
A further limitation to our “gaming” experiment (and all manipulation experiments) is that they produce outputs which may be viewed as “out of distribution”; but generally the purpose of factuality metrics is to be able to apply them to arbitrary model outputs and corresponding references, so it is not entirely clear what “in distribution” means under this assumption.
We also note that we primarily stress-test metrics designed for English. The impact of our translated gamed phrases and manipulations on multilingual metrics remains unclear.
Finally, while we stress-test metrics, we do not propose a solution to mitigate their flaws. However, similar to prior work gabriel-etal-2021-go; chen-etal-2021-factuality-checkers, we highlight specific issues in current metrics such as reliance on superficial cues, sensitivity to benign changes over inaccuracies, and score inflation from innocuous assertions, aiming to inform the development of more reliable metrics.
Despite these limitations, we believe that—taken together—our results suggest that one should be careful about their interpretation of automatic factual consistency metrics.
Acknowledgements
This work was supported in part by the National Science Foundation (NSF), grant 2211954. This work was also supported by the Wellcome Trust, grant 313618/Z/24/Z.
Appendix
8 Benchmark Datasets
We list the datasets used in the benchmarks and specify which split (train/test) they are included in Table 3
@@ 8 Benchmark Datasets
LLM-Aggrefact [tang2024minicheck] TofuEval-MeetB [tang-etal-2024-tofueval], ExpertQA [malaviya2023expertqa], Lfqa [chen2023understanding], RAGTruth [wu2023ragtruth], FactCheck-GPT [wang2023factcheck], Wice [kamoi-etal-2023-wice], TofuEval-MediaS [tang-etal-2024-tofueval], ClaimVerify [liu-etal-2023-evaluating], Reveal [jacovi2024chain]
Aggrefact [tang2022understanding] FRANK [pagnoni2021understanding], Polytope [huang-etal-2020-achieved], FactCC [kryscinski-etal-2020-evaluating], Goyal21 [goyal-durrett-2021-annotating], Wang20 [wang2020asking], Cao22 [cao-etal-2022-hallucinated], XSumFaith [maynez-etal-2020-faithfulness], CLIFF [cao-wang-2021-cliff], SummEval [fabbri2021summeval]
FacEval [wang2022analyzing] SAMSUM [gliwa2019samsum]
LLM-Dialogue [ramprasad-etal-2024-analyzing] SAMSUM [gliwa2019samsum], DialogSum [chen-etal-2021-dialogsum]
Genaudit [krishna2024genaudit] XSUM [narayan-etal-2018-dont] REDDIT, ACIBENCH [yim2023aci]
Table 3: Datasets included in the train and test splits for each benchmark.
9 Heuristic Indicators of “Factuality”
Another predictor we use is semantic similarity. Specifically, embed summary and source sentences via BERT [devlin2018bert], then score each summary sentence based on which source sentence it is most similar to (using cosine similarity). We average all summary sentence scores to derive a final score.
10 Shallow MLP Model
We train an MLP classification model to predict factuality labels using shallow features, with two hidden layers of dimensions 100 and 50 with a learning rate of 0.001.
11 Benign Summary Manipulations
Table 4 lists the prompts used with GPT-40-mini to generate summary variants, each designed to vary in specific ways without affecting factual consistency.
We spot-check examples to ensure that manipulations preserve factual meaning and do not introduce contradictions or new factual errors. However, this process is not exhaustive, and undetected issues may introduce some noise into our results.